OpenStudy (anonymous):

solve system of equations by substitution method -x+2y=3 x^2+y=3

OpenStudy (mrnood):

Re-arrange the 2nd equation so it reads y=......

OpenStudy (john_es):

Or you can start from the first equation, \[y=(3+x)/2\] And then substitute in the second equation.

OpenStudy (mrnood):

then in the first equation where there is y substitute with the expression you derived above. That will give you an equation in x only - so you can solve for x

OpenStudy (anonymous):

i got stuck at -2x^2-x+9=0

OpenStudy (anonymous):

did i make a mistake or does it have 2 x values?

OpenStudy (john_es):

The equation should be, \[x^2+\frac{3+x}{2}=3\Rightarrow 2x^2+x-3\] And then solve the second grade equation.

OpenStudy (john_es):

I mean, \[2x^2+x−3=0\]

OpenStudy (anonymous):

how would i deal with the 2x^2?

OpenStudy (john_es):

With the formula, \[ax^2+bx+c=0\Rightarrow x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

Ohhhhhhhh i see. thanks for all the help!! i think i got it :)))

OpenStudy (john_es):

Ok, ;).